A coil of inductance 2mH and resistance 15Ω is connected with in parallel ith capacitor of 0.01μF. The peak oscillator current at resonance is?
1
Expert's answer
2020-06-17T09:29:02-0400



As per phaser diagram, For circuit to be resonance, Impedance (resistance) of circuit must be minimum.


So,

"I_C = I_Lsin\theta"


simply putting value of "I_C , I_L" and "sin\theta"


"\frac{V}{X_C} = \frac{V}{Z_L}(\frac{X_L}{Z_L})"


"X_CX_L = Z_L^2" . . . . . . (i)


since, "X_C = \frac{1}{\omega C}" , "X_L = {\omega L}" and "Z_L^2 = R^2 + X_L^2"


then equation (i) will be,


"\frac{\omega L}{\omega C} = R^2 + (\omega L)^2"


"(\omega)^2 = (2\pi f_r)^2= [\frac{1}{LC} - {\frac{R}{L^2}}]"


"f_r =(\frac{1}{2\pi}) \sqrt{\frac{1}{LC} - {\frac{R}{L^2}}}" is the resonance frequency.


Again,

"\frac {V}{Z_f} = \frac{V}{Z_L}(\frac{R}{Z_L})" where "Z_r" is the total effective impedance of the circuit at resonance.


"Z_f = {\frac{Z_L^2}{R}}"

using (i) again here,


"Z_f = \frac{X_CX_L}{R} = \frac{\omega L}{R \omega C} = \frac{L}{RC}"


Now using these formulas,

Impedance of the circuit will be,

"Z_f = \frac{L}{LC} = \frac{2*10^-3}{15*0.01*10^-6} = 13.33*10^3 \Omega"



For current, use ohm's law,


"I = \frac{V}{Z_f} = \frac{VCR}{L}" . . . . . . . (ii)



If V = 1V then current will be, using equation(ii) putting value of "Z_f"


"I = 0.075*10^{-3} A"